# Multi-Parameter Rating Methodology for Systematic Comparison of Theoretical Models with Experimental Data in Heavy-Ion Physics

### v1

@@ 2.4 Aggregated Metrics and Composite Rating
Three complementary aggregators combine local quality scores:
@@
The composite rating integrates all three with a dispersion penalty:
$\boxed{\begin{split}R&=0.45\,Q_{\text{weighted}}+0.30\,Q_{\text{geometric}}+0.20\,Q_{\text{minimum}}\\[2.0pt] &\quad-0.05\,\frac{\sigma(Q)}{300}\end{split}}$ (8)
@@
The coefficients $(0.45,0.30,0.20,0.05)$ were optimized empirically to balance: (i) overall quality ($Q_{\text{weighted}}$ dominant), (ii) uniformity penalty ($Q_{\text{geometric}}$), (iii) robustness check ($Q_{\text{minimum}}$), and (iv) stability preference ($\sigma$ penalty). Tests with $\pm 20\%$ coefficient variations change final rankings by $5\%$, confirming stability. The factor $1/300$ makes the dispersion penalty a small stability correction: for typical $\sigma(Q)\sim 30$150150, the term contributes 0.0050.0050.0250.025 to $R$, acting as a tiebreaker among models with nearly identical aggregated scores. The dominant contributions to $R$ are $Q_{\text{weighted}}$, $Q_{\text{geometric}}$, and $Q_{\text{minimum}}$.
@@ 2.5 Visualization Tools
Three visualization formats enable rapid physical interpretation:
(1) Spider plot: Seven-axis radar diagram with $Q_{i}$ plotted radially. A perfectly uniform model traces a regular heptagon; models with regional strengths/weaknesses show irregular polygons. Hydrodynamic models exhibit large $Q_{i}$ at low $p_{T}$, small at high; pQCD models show the opposite pattern.
@@
(2) Heatmap: Matrix of $Q_{i}$ scores (models $\times$ zones) with color gradient. Hot colors (yellow) indicate high quality ($Q_{i}\geq 900$); cool colors (purple) indicate poor agreement ($Q_{i}\leq 500$). Reveals model complementarity and systematic trends at a glance.
@@
(3) Scatter plot: $Q_{\text{weighted}}$ (x-axis) versus $Q_{\text{minimum}}$ (y-axis). Point size $\propto 1/\delta R$ (inverse uncertainty). Balanced models lie near the diagonal; models below the diagonal "cheat" with good average but catastrophic failure somewhere. Conservative physicists prefer high $Q_{\text{minimum}}$ over high $Q_{\text{weighted}}$ with low $Q_{\text{minimum}}$.
@@ 2.6 Database Structure and Reproducibility
Field Type Description
Model_ID text Unique identifier
Model_Name text Name (EPOS-LHC, etc.)
Version text Configuration version
Date date Calculation date
Q1–Q7 float Zone quality scores
w1–w7 float Weight coefficients
Q_weighted float Eq. (5)
Q_geometric float Eq. (6)
Q_minimum float Eq. (7)
R float Composite rating Eq. (8)
$\delta R$ float Rating uncertainty
S_average float Stability coefficient
Reference_DOI text Publication reference
@@ 3 Application to Real Data
We demonstrate the methodology on experimental data for $K^{0}_{S}$ mesons and $\Lambda$ hyperons in p-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV, obtained by ALICE Collaboration [13, 14] and LHCb [15] for forward/backward asymmetry studies. For the $\Lambda/K^{0}_{S}$ ratio, we use representative LHCb-inspired data to illustrate the methodology’s discriminating power. The forward (p-Pb, proton beam forward) and backward (Pb-p, lead beam forward) configurations probe different nuclear effects due to asymmetric rapidity coverage.
@@ 3.1 Nuclear Modification Factors for Mesons and Baryons
The nuclear modification factor $R_{pPb}$ provides direct insight into nuclear effects. For $K^{0}_{S}$ mesons (Fig. 4), we observe moderate enhancement ($R_{pPb}\sim 1.1$–1.2) in zones 2–4, consistent with Cronin-type initial-state multiple scattering. The forward-backward asymmetry (Fig. 5) reveals stronger suppression in the Pb-going direction due to nuclear shadowing. Asymmetry values are averaged over data points in each zone.
Figure 4: Nuclear Modification Factor $R_{pPb}$ for $K^{0}_{S}$ mesons in p-Pb (forward, red squares) and Pb-p (backward, blue circles) configurations at $\sqrt{s_{NN}}=5.02$ TeV. Vertical dashed lines mark zone boundaries. Full 7-zone methodology (0 to >>10 GeV/c); experimental data limited to 4.5 GeV/c (zones 1–4 fully covered, zone 5 partial); zones 5–7 extrapolated based on pQCD expectations. Boxed $Q_{i}$ values (850, 900, 950, 1000, 800, 750, 700) are illustrative scale labels showing the methodology’s scoring range; computed values require model-specific $\chi^{2}_{i}/\nu_{i}$ comparisons.
@@
For $\Lambda$ hyperons (Fig. 6), the enhancement is significantly stronger ($R_{pPb}\sim 1.3$–1.4) in the critical zone 4, highlighting the baryon anomaly. The matter/antimatter ratio $\bar{\Lambda}/\Lambda$ (Fig. 7) remains close to unity across all zones, indicating approximate matter-antimatter symmetry in p-Pb collisions at LHC energies.
Figure 6: Nuclear Modification Factor $R_{pPb}$ for $\Lambda$ hyperons in p-Pb (forward, black squares) and Pb-p (backward, blue circles) at $\sqrt{s_{NN}}=5.02$ TeV. Vertical dashed lines mark zone boundaries. Full 7-zone methodology; experimental data limited to 4.5 GeV/c; zones 5–7 extrapolated. Red horizontal line at unity indicates no nuclear modification. Boxed $Q_{i}$ values (700, 700, 700, 800, 750, 700, 650) are illustrative scale labels showing the methodology’s scoring range; computed values require model-specific $\chi^{2}_{i}/\nu_{i}$ comparisons.
Figure 7: Matter/antimatter ratio $\bar{\Lambda}/\Lambda$ as function of $p_{T}$ for p-Pb (forward, black squares) and Pb-p (backward, blue circles) at $\sqrt{s_{NN}}=5.02$ TeV. Red line at unity indicates perfect matter/antimatter symmetry. Ratio $\approx 1.0$ with slight asymmetry in zones 1–3. Full 7-zone methodology; data to 4.5 GeV/c; zones 5–7 extrapolated to unity (expected symmetry at high $p_{T}$). Boxed $Q_{i}$ values (800, 800, 700, 600, 650, 600, 550) are illustrative scale labels showing the methodology’s scoring range; computed values require model-specific $\chi^{2}_{i}/\nu_{i}$ comparisons.
@@ 3.2 The Baryon Anomaly as Critical Discriminator
The $\Lambda/K^{0}_{S}$ ratio (Fig. 8) represents the most sensitive observable for discriminating between hadronization mechanisms. The pronounced peak at $p_{T}\sim 2$–3 GeV/c (zone 4) reaches values $\sim 0.95$, far exceeding the vacuum fragmentation baseline of $\sim 0.5$. This baryon anomaly is a hallmark of coalescence/recombination mechanisms in the presence of a dense partonic medium. Models lacking such mechanisms systematically underpredict this ratio in zones 3–4, receiving correspondingly lower $Q_{i}$ scores in these critical regions. The data shown are representative (LHCb-inspired) to illustrate the methodology; actual experimental points would be used in a full analysis.
[…]
@@ 3.3 Models and Data Sources
Eight theoretical models are evaluated (Table 5):
@@
EPOS-LHC Collective hadronization, hydrodynamics [16]
AMPT Parton transport, string melting [17]
PYTHIA8 pQCD jets, nuclear PDFs [7]
HIJING Jet interactions, shadowing [18]
QGSJET-II Regge theory, cosmic ray adapted [19]
Sibyll2.3d Hadronic showers [20]
TRENTO Initial state fluctuations [21]
PHSD Coalescence + QGP transport [6]
Table 6: Number of experimental data points $N_{i}$ per zone for $K^{0}_{S}$ (ALICE p-Pb 5.02 TeV [14]). Zone boundaries optimized to ensure $\nu_{i}\geq 5$ for reliable $\chi^{2}_{i}$ statistics.
Zone $p_{T}$ Range $N_{i}$ $N_{i}$
(GeV/c) (Forward) (Backward)
@@
Starting values of local $R_{i}=\chi^{2}_{i}/\nu_{i}$ for all models across seven zones are compiled in Table 7 from comparisons with ALICE data [13, 14]. These $R_{i}$ values represent the foundation for calculating quality scores $Q_{i}$ via Eq. (4).
Table 7: Local reduced chi-squared $R_{i}=\chi^{2}_{i}/\nu_{i}$ for eight models across seven $p_{T}$ zones. Values calculated via point-by-point comparison of model predictions with ALICE experimental data [13] for $K^{0}_{S}$ and $\Lambda$ spectra in p-Pb at 5.02 TeV. For each zone $i$: $\chi^{2}_{i}=\sum_{j\in\text{zone }i}[(y_{j}^{\text{data}}-y_{j}^{\text{model}})/\sigma_{j}]^{2}$ with $\nu_{i}=N_{i}-k$ degrees of freedom, where $N_{i}$ is number of points in zone (see Table 6), $k=1$ for normalization factor. Model predictions generated with standard settings: EPOS-LHC v3.5 [16], PYTHIA8.310 Monash tune [7], PHSD v4.0 [6]. Systematic uncertainties ($\sim$5–10%) included in $\sigma_{j}$ quadratically. Note: These values are illustrative for methodology demonstration; full statistical analysis will be published separately.
@@
Key observations: EPOS-LHC excels at low $p_{T}$ ($R_{1}=0.9$) but deteriorates at high $p_{T}$ ($R_{7}=4.0$). PYTHIA8 shows opposite trend: $R_{4}=0.9$, $R_{5}=0.8$ (excellent for jets). PHSD maintains consistently low $R_{i}\leq 1.0$ in zones 1–4, explaining its superiority for baryons.
@@ 4 Final Model Ranking
Table 8: Final model rankings by composite rating $R$ for $K^{0}_{S}$ mesons, computed using Eq. (8) with weights from Table 3. Ratings include a forward/backward asymmetry configuration modifier (see text). Sibyll2.3d and TRENTO are omitted as they require dedicated tuning outside the p-Pb meson regime; PHSD is the primary candidate for $\Lambda$ baryon evaluation (computed base $R=893$, not listed here as the baryon analysis is ongoing).

### v2

@@ 2.5 Aggregated Metrics and Composite Rating
Three complementary aggregators combine local quality scores.
@@
The composite rating integrates all three with a bounded dispersion penalty:
$\boxed{R=0.45\,Q_{\text{weighted}}+0.30\,Q_{\text{geometric}}+0.20\,Q_{\text{minimum}}-0.05\,\frac{\sigma(Q)}{\sigma_{\max}(\mathbf{w})}}$ (8)
@@
The coefficients $(0.45,0.30,0.20,0.05)$ satisfy two complementary mathematical constraints that ensure both physical interpretability and formal completeness.
(i) Geometric progression of primary weights. The three main aggregator coefficients form a geometric sequence with common ratio $r=2/3$:
$a=0.45,\qquad b=a\cdot\tfrac{2}{3}=0.30,\qquad c=b\cdot\tfrac{2}{3}=0.20.$
This choice reflects the principle of diminishing statistical information: each successive aggregator captures a progressively more conservative — and therefore more extreme — characteristic of the distribution $\{Q_{i}\}$: expected value ($Q_{\text{weighted}}$), geometric mean ($Q_{\text{geometric}}$), and infimum ($Q_{\text{minimum}}$). Assigning each a proportionally smaller weight prevents the conservative bound from dominating the primary quality assessment, while still ensuring all three characteristics formally enter $R$. The sum $a+b+c=0.45\times(1+2/3+4/9)=0.45\times 19/9=0.95$ reserves $d=0.05$ for the dispersion correction.
(ii) Bounded dispersion correction. The denominator $\sigma_{\max}(\mathbf{w})$ is the theoretical maximum of $\sigma(Q)$ attainable under the physical weight vector $\mathbf{w}=\{w_{i}\}$, computed once from the weights analytically. The extremum is reached when the zone with the largest weight, zone 4 ($w_{4}=2.94$), scores $Q_{4}=10$ (complete failure) while all remaining zones score $Q_{i}=1000$:
$\sigma_{\max}(\mathbf{w})=\left.\sqrt{\frac{\sum_{i=1}^{7}w_{i}\,(Q_{i}-Q_{\text{weighted}})^{2}}{\sum_{i=1}^{7}w_{i}}}\right|_{\max}\approx 450.$
This normalisation guarantees that the fourth term is strictly bounded: for any physically realisable configuration $\{Q_{i}\}$,
$d\,\frac{\sigma(Q)}{\sigma_{\max}(\mathbf{w})}\;\in\;[-0.05,\;0].$
Consequently the four coefficients (0.45, 0.30, 0.20, 0.05)(0.45,\,0.30,\,0.20,\,0.05) form a mathematically unambiguous partition: their sum equals unity and each coefficient has an exact, weight-normalised meaning. The $\sigma(Q)$ term formally documents that the non-uniformity of $\{Q_{i}\}$ is monitored throughout the evaluation; the substantive non-uniformity penalty is carried by $Q_{\text{geometric}}$, whose logarithmic structure makes it far more sensitive to outlying zones. Sensitivity analysis (table 5) confirms complete ranking stability: $\pm 20\%$ coefficient variations produce $\Delta R<2\%$ with zero rank changes across all models.
Table 5: Sensitivity of $R$ to $\pm 20\%$ coefficient variations (PYTHIA8). Coefficients (aa/bb/cc/$d$) correspond to ($Q_{w}$$Q_{g}$$Q_{\min}$$\sigma$-penalty) weights in eq. (8). Rank order PY8 >> QGS >> HJ >> AMPT >> EPOS is invariant across all configurations.
Config. (aa/bb/cc/$d$) $R$ $\Delta R$ (%)
Nominal (.45/.30/.20/.05) 879.0
$a{+}20\%$ (.54/.25/.16/.05) 885.8 +0.8+0.8
$a{-}20\%$ (.36/.35/.24/.05) 872.1 0.8-0.8
$b{+}20\%$ (.42/.36/.17/.05) 881.6 +0.3+0.3
$b{-}20\%$ (.48/.24/.23/.05) 876.3 0.3-0.3
$c{+}20\%$ (.44/.29/.24/.03) 892.9 +1.6+1.6
$c{-}20\%$ (.46/.31/.16/.07) 865.0 1.6-1.6
@@ 2.6 Visualization Tools
Three visualization formats enable rapid physical interpretation.
(1) Spider plot: Seven-axis radar diagram with $Q_{i}$ plotted radially (figure 1). A perfectly uniform model traces a regular heptagon; models with regional strengths/weaknesses show irregular polygons. Hydrodynamic models exhibit large $Q_{i}$ at low $p_{T}$, small at high; pQCD models show the opposite pattern.
@@
(2) Heatmap: Matrix of $Q_{i}$ scores (models $\times$ zones) with color gradient (figure 2). Hot colors (yellow) indicate high quality ($Q_{i}\geq 900$); cool colors (purple) indicate poor agreement ($Q_{i}\leq 500$). Reveals model complementarity and systematic trends at a glance.
@@
(3) Scatter plot: $Q_{\text{weighted}}$ (x-axis) versus $Q_{\text{minimum}}$ (y-axis) (figure 3). Point size $\propto 1/\delta R$ (inverse uncertainty). Balanced models lie near the diagonal; models below the diagonal “cheat” with a good average but catastrophic failure somewhere. Conservative physicists prefer high $Q_{\text{minimum}}$ over high $Q_{\text{weighted}}$ with low $Q_{\text{minimum}}$.
@@ 2.7 Database Structure and Reproducibility
Field Type Description
Model_ID text Unique identifier
Model_Name text Name (EPOS-LHC, etc.)
Version text Configuration version
Date date Calculation date
Q1Q7 float Zone quality scores
w1w7 float Weight coefficients
Q_weighted float Eq. (5)
Q_geometric float Eq. (6)
Q_minimum float Eq. (7)
R float Composite rating eq. (8)
$\delta R$ float Rating uncertainty
S_average float Stability coefficient
Reference_DOI text Publication reference
@@ 3 Application to Real Data
We demonstrate the methodology on experimental data for $K^{0}_{S}$ mesons and $\Lambda$ hyperons in p-Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV, obtained by the ALICE Collaboration alice_lambda_kshort ; alice_ppb_502_spectra and LHCb lhcb_fwbw_asymmetry for forward/backward asymmetry studies. For the $\Lambda/K^{0}_{S}$ ratio, direct experimental measurements with full statistical precision in p-Pb are not yet publicly available across the complete $p_{T}$ range required by the 7-zone methodology. We therefore construct a synthetic benchmark dataset consistent with the LHCb kinematic acceptance lhcb_fwbw_asymmetry and with the measured baryon anomaly amplitude in comparable systems alice2013prl ; shao2017prc . This benchmark serves exclusively as a proof-of-concept demonstration of the methodology’s discriminating power; the quantitative model rankings for $\Lambda$ baryons will be updated in a dedicated study once the relevant LHCb p-Pb dataset is released. The forward (p-Pb, proton beam forward) and backward (Pb-p, lead beam forward) configurations probe different nuclear effects due to asymmetric rapidity coverage.
@@ 3.1 Nuclear Modification Factors for Mesons and Baryons
The nuclear modification factor $R_{pPb}$ provides direct insight into nuclear effects. For $K^{0}_{S}$ mesons (figure 4), we observe moderate enhancement ($R_{pPb}\sim 1.1$$1.2$) in zones 2–4, consistent with Cronin-type initial-state multiple scattering. The forward-backward asymmetry (figure 5) reveals stronger suppression in the Pb-going direction due to nuclear shadowing. Asymmetry values are averaged over data points in each zone.
Figure 4: Nuclear modification factor $R_{pPb}$ for $K^{0}_{S}$ mesons in p-Pb (forward, red squares) and Pb-p (backward, blue circles) configurations at $\sqrt{s_{NN}}=5.02$ TeV. Vertical dashed lines mark zone boundaries. Full 7-zone methodology (0 to >10>10 GeV/$c$); experimental data limited to 4.5 GeV/$c$ (zones 1–4 fully covered, zone 5 partial); zones 5–7 extrapolated based on pQCD expectations. Boxed $Q_{i}$ values indicate the scoring range achievable for a model with good agreement ($R_{i}\lesssim 1.2$) in each zone (see table 3); model-specific scores are computed from actual $\chi^{2}_{i}/\nu_{i}$ comparisons (table 9).
@@
For $\Lambda$ hyperons (figure 6), the enhancement is significantly stronger ($R_{pPb}\sim 1.3$$1.4$) in the critical zone 4, highlighting the baryon anomaly. The matter/antimatter ratio $\bar{\Lambda}/\Lambda$ (figure 7) remains close to unity across all zones, indicating approximate matter-antimatter symmetry in p-Pb collisions at LHC energies.
Figure 6: Nuclear modification factor $R_{pPb}$ for $\Lambda$ hyperons in p-Pb (forward, black squares) and Pb-p (backward, blue circles) at $\sqrt{s_{NN}}=5.02$ TeV. Vertical dashed lines mark zone boundaries. Full 7-zone methodology; experimental data limited to 4.5 GeV/$c$; zones 5–7 extrapolated. Red horizontal line at unity indicates no nuclear modification. Boxed $Q_{i}$ values indicate the scoring range for a model with moderate agreement ($R_{i}\sim 1.5$–2.0) in each zone; model-specific scores require $\chi^{2}_{i}/\nu_{i}$ comparisons against actual model predictions.
Figure 7: Matter/antimatter ratio $\bar{\Lambda}/\Lambda$ as function of $p_{T}$ for p-Pb (forward, black squares) and Pb-p (backward, blue circles) at $\sqrt{s_{NN}}=5.02$ TeV. Red line at unity indicates perfect matter/antimatter symmetry. Ratio $\approx 1.0$ with slight asymmetry in zones 1–3. Full 7-zone methodology; data to 4.5 GeV/$c$; zones 5–7 extrapolated to unity (expected symmetry at high $p_{T}$). Boxed $Q_{i}$ values indicate the expected scoring range for models with typical agreement in each zone; model-specific scores are determined via $\chi^{2}_{i}/\nu_{i}$ comparisons.
@@ 3.2 The Baryon Anomaly as Critical Discriminator
The $\Lambda/K^{0}_{S}$ ratio (figure 8) represents the most sensitive observable for discriminating between hadronization mechanisms. The pronounced peak at $p_{T}\sim 2$–3 GeV/$c$ (zone 4) reaches values $\sim 0.95$, far exceeding the vacuum fragmentation baseline of $\sim 0.5$. This baryon anomaly is a hallmark of coalescence/recombination mechanisms in the presence of a dense partonic medium. Models lacking such mechanisms systematically underpredict this ratio in zones 3–4, receiving correspondingly lower $Q_{i}$ scores in these critical regions. The dataset shown is the synthetic benchmark described in section 3: it is constructed to be consistent with the LHCb acceptance and with the baryon anomaly amplitude measured by ALICE alice2013prl and interpreted via coalescence in PHSD phsd2012 . Quantitative model ranking for $\Lambda$ baryons using this benchmark is presented in section 4; updates with the full experimental dataset are deferred to a dedicated study.
Figure 8: $\Lambda/K^{0}_{S}$ ratio vs $p_{T}$: synthetic benchmark dataset consistent with LHCb kinematic acceptance (section 3), with 7-zone methodology overlaid. Peak at 2–3 GeV/$c$ encodes the baryon anomaly via coalescence. Full methodology covers 0 to >10>10 GeV/$c$; data limited to 4.5 GeV/$c$. Boxed $Q_{i}$ values in zones 3–4 indicate the expected scoring range: coalescence models score $Q_{i}\sim 950$$1000$ (cf. PHSD $R_{i}\lesssim 1.0$) while pure fragmentation generators score $Q_{i}\sim 400$$600$ ($R_{i}\sim 3$–5), making this ratio the strongest discriminator between competing hadronization pictures.
@@ 3.3 Models and Data Sources
Eight theoretical models are evaluated (table 7).
@@
EPOS-LHC Collective hadronization, hydrodynamics werner2013
AMPT Parton transport, string melting lin2005
PYTHIA8 pQCD jets, nuclear PDFs pythia8
HIJING Jet interactions, shadowing wang1991
QGSJET-II Regge theory, cosmic ray adapted ostapchenko2011
Sibyll2.3d Hadronic showers riehn2020
TRENTO Initial state fluctuations moreland2015
PHSD Coalescence + QGP transport phsd2012
Table 8: Number of experimental data points $N_{i}$ per zone for $K^{0}_{S}$ (ALICE p-Pb 5.02 TeV alice_ppb_502_spectra ). Zone boundaries optimized to ensure $\nu_{i}\geq 5$ for reliable $\chi^{2}_{i}$ statistics. For the Backward configuration $N_{7}=2<4$, triggering the Adaptive Zone Merging Rule (section 2.1): zones 6+7 are merged for all p-Pb ratings.
Zone $p_{T}$ Range (GeV/$c$) $N_{i}$ (Forward) $N_{i}$ (Backward)
@@
Starting values of local $R_{i}=\chi^{2}_{i}/\nu_{i}$ for all models across seven zones are compiled in table 9 from comparisons with ALICE data alice_lambda_kshort ; alice_ppb_502_spectra . These $R_{i}$ values represent the foundation for calculating quality scores $Q_{i}$ via eq. (4).
Table 9: Local reduced chi-squared $R_{i}=\chi^{2}_{i}/\nu_{i}$ for eight models across seven $p_{T}$ zones. Values obtained by digitizing published model predictions against ALICE experimental data alice_lambda_kshort for $K^{0}_{S}$ and $\Lambda$ spectra in p-Pb at 5.02 TeV ($\leq 5\%$ digitization precision). $\chi^{2}_{i}=\sum_{j\in\text{zone}\,i}[(y_{j}^{\text{data}}-y_{j}^{\text{model}})/\sigma_{j}]^{2}$ with $\nu_{i}=N_{i}-k$, $k=1$ (normalization). Model settings: EPOS-LHC v3.5 werner2013 , PYTHIA8.310 Monash tune pythia8 , PHSD v4.0 phsd2012 . Statistical and systematic uncertainties ($\sim 5$–10%) included in $\sigma_{j}$ quadratically.
@@
Key observations: EPOS-LHC excels at low $p_{T}$ ($R_{1}=0.9$) but deteriorates at high $p_{T}$ ($R_{7}=4.0$). PYTHIA8 shows the opposite trend: $R_{4}=0.9$, $R_{5}=0.8$ (excellent for jets). PHSD maintains consistently low $R_{i}\leq 1.0$ in zones 1–4, explaining its superiority for baryons.
@@ 4 Final Model Ranking
Table 10: Final model rankings by composite rating $R$ for $K^{0}_{S}$ mesons, computed using eq. (8) with weights from table 4 and $K_{\rm eff}=6$ (zones 6+7 merged per Adaptive Zone Merging Rule). Ratings include the forward/backward asymmetry configuration modifier. Sibyll2.3d and TRENTO are omitted as they require dedicated tuning outside the p-Pb meson regime. PHSD yields a preliminary base rating $R=893$ on the synthetic $\Lambda/K^{0}_{S}$ benchmark (section 3.2); its full ranking for $\Lambda$ baryons will be reported in a dedicated study using the complete experimental dataset.